OpenStudy (selena2345):

Evaulate The Expression a-b if a =4 and b=12

OpenStudy (anonymous):

so then you would substitute those values into the equation since a =4 and b = 12 a-b = 4-12 = -8

OpenStudy (selena2345):

Thank You Alexeis_Nicole

OpenStudy (anonymous):

yep np. so how about if a=6 and b=8 and it asked you to evaluate for a-b what would you do?

OpenStudy (selena2345):

6-8=-2

OpenStudy (anonymous):

yeep :) goodjob!

OpenStudy (anonymous):

Sorry I'm not so good at math myself lol